### v1

@@ 3.2 Estimating network parameters
Our qualitative analysis showed that estimators (21) and (20), while analytically interesting, are subpar to the estimator from Formula (19). As such we omit them from further study.
3.3 Results
@@
Figures 2 and 3 present the observed and modelled in-degree densities in the communities of the real networks. We estimated the densities from our predicted degree sequence for $\rho_{i}$ values obtained using estimator (19). First, looking at Figure 3, we observe that all our modelled densities fit well to the observed degree sequences. Our model fits especially well to the tails of the densities while keeping the peak well approximated in the most clusters as well. The other $\rho_{i}$ estimators often fitted better to the peaks but produced unacceptable estimation of the tails.
@@
Similarly, as per Figure 2, our modelled densities fits well to most of the clusters with again an especially good fit to the tails while keeping the peaks closely approximated as well. Notably, there are clusters with degree sequences that our model was not able to capture properly, but we believe that this is to be expected from real dataset where the behaviour of the network is not fully predictable. We should also note that our parameter estimates and degree estimates are all related to each other through the equations. That is, we cannot improve a fit to one community (e.g., “Control theory”) without risking that fit to another community (e.g., ”Computer science”) suffers a decline in quality. As such, the quality of the density estimation should be considered in total for all communities simultaneously. As for fitting to the whole network, we observe that the density resulting from the standard 3DSI model is almost identical to the density from our model aside from the tail, which suggests that we can use the standard model to approximate the global degree sequence. The parameters for the standard model have been estimated as weighted averages over the communities with probabilities $p_{i}$ as weights. Overall, the results indicate that our model fits well to real network data, with the Gini-based $\rho_{i}$ estimation producing the high quality results with good tail estimates. As such, our model offers a unique possibility to analytically analyse the in-degree sequences of any network with the in-degrees of each community considered separately.
@@
Table 2 gives the estimated values of the parameters of our model for both networks. We observe that there is a large variance of $\hat{\rho}$ in a single network which confirms that different disciplines tend to cite papers with different ratios of accidentality-to-preferentiality.
@@ 4 Conclusion
To summarise, we have devised an analytical in-degree sequence model for citation networks that is a direct generalisation of the 3DSI model but accounts for the community structure in the graph. We note that it exactly mimics the original 3DSI model with no communities ($k=1$) assuming no loops are allowed. Overall, our model has nice analytical properties and allows for the derivation of many exact and asymptotic results, as well as fits well to some real networks. Since it outputs a degree sequence for each community separately, it offers unique insight not seen previously in citation network research.
We will focus our future works on devising an algorithmic framework for generating benchmark graphs that can be used for community detection, as we believe that our model will be complementary to approaches like SBM or LFR in that it allows for flexible modification of the in-degree sequences inside the communities, allowing for modelling broad families of networks. It would also be beneficial to consider further generalizations of our model in which accidental edges are also distributed non-homogenously, or in which the parameter values depend on the size of the network – recent research indicates that interdisciplinary citations trends change non-randomly in time Rong et al. (2025).

### v2

@@ 3.2 Estimating network parameters
Our qualitative analysis showed that estimators (20) and (21), while analytically interesting, are subpar to the estimator from Formula (19). As such we omit them from further study. However, in Appendix A, we will also consider numerical minimisation of the mean squared log loss.
3.3 In-degree fit with exogenous communities
@@
Figures 2 and 3 present the observed and modelled complementary cumulative distribution functions in the communities of the real networks. We estimated all values from our predicted degree sequence for $\rho_{i}$ values obtained using estimator (19). First, looking at Figure 3, we observe that all our modelled densities fit well to the observed degree sequences. Our model fits especially well when describing both low- and medium-degree nodes and only overestimated largest in-degree in biggest communities. The other $\rho_{i}$ estimators often fitted better near distribution means but diverged from real distributions at tails.
@@
Similarly, as per Figure 2, our modelled densities fits well to most of the clusters with again an especially good fit to low- and medium-degree nodes while keeping the highest degrees bounded from above. Notably, the fit is extremely good for all but the two biggest communities. There are clusters with degree sequences that our model was not able to capture properly, but we believe that this is to be expected from real dataset where the behaviour of the network is not fully predictable. We should also note that our parameter estimates and degree estimates are all related to each other through the equations. That is, we cannot improve a fit to one community (e.g., “Control theory”) without risking that fit to another community (e.g., ”Computer science”) suffers a decline in quality. As such, the quality of the density estimation should be considered in total for all communities simultaneously. As for fitting to the whole network, we observe that the distribution resulting from the standard 3DSI model is almost identical to the distribution from our model aside from the tail, which suggests that we can use the standard model to approximate the global degree sequence. The parameters for the standard model have been estimated as weighted averages over the communities with probabilities $p_{i}$ as weights. Overall, the results indicate that our model fits well to real network data, with the Gini-based $\rho_{i}$ estimation producing the high quality results with good tail estimates. The biggest discrepancy happens in high-degree nodes, which our model consistently overestimates. This effect may have many explanations, but in the context of citation networks the occurrence of ageing or attention decay seems the most likely. As our model does not take the attention decay into account, the overestimation is expected for large communities. In Appendix A, we propose an alternative version of our model that aims to alleviate this issue. Nonetheless, the problem does not occur for the remaining clusters.
Overall, our model offers a unique possibility to analytically analyse the in-degree sequences of any network with the in-degrees of each community considered separately.
@@
Table 2 gives the estimated values of the parameters of our model for both networks. We observe that there is a large variance of $\hat{\rho}$ in a single network which confirms that different disciplines tend to cite papers with different ratios of accidentality-to-preferentiality. Our results regarding the Cora network confirm the observations made in Mrowinski et al. (2022), where the authors showed that theoretical fields tend to be governed in larger part by accidental citations, while the applied fields have higher preferentiality. This effect is not clearly visible in our results on the DBLP authors dataset, which is likely caused by either different type of topology or large inequality in cluster sizes.
@@
3.4 In-degree fit with communities identified by the Leiden algorithm
To confirm that our model works not only with exogenous labels, we repeat the comparison on the Cora dataset but instead of using the ground-truth label vector, we first run the Leiden Traag et al. (2019) algorithm to find the communities. The algorithm is run from a random initial membership until no improvement in modularity is detected. The CCDF comparison on the seven largest found communities, as well as the full network, can be seen in Figure 4. We selected seven largest communities because the Leiden algorithm found many (100+) clusters, for many of which fitting our model would be pointless due to low sample size and high noise.
Similarly to what we have seen on real data, we observe a good fit to almost all large communities, both in peaks and tails. The ageing bias is not visible likely due to the fairly small sizes of all communities. Overall, we found no indicator that our model should be limited to exogenous communities and seems to work correctly also with algorithmically-detected associative communities. One note to be made is that for extremely small communities, our model cannot be reasonably fitted, and the in-degrees corresponding to such nodes were omitted from the full network CCDF plot.
Figure 4: Complementary cumulative distribution functions of the observed and modelled in-degree sequences in the seven largest clusters of the Cora dataset with Leiden-assigned labels. Our model was fitted using Gini-based $\rho_{i}$ estimators. The last plot shows the full in-degree sequences.
@@ 4 Conclusion
We have devised an analytical in-degree sequence model for citation networks that is a direct generalisation of the 3DSI model accounting for the community structure in the graph. We note that it exactly mimics the original 3DSI model with no communities ($k=1$) assuming no loops are allowed. Overall, our model has nice analytical properties and allows for the derivation of many exact and asymptotic results, as well as fits well to some real networks. Since it outputs a degree sequence for each community separately, it offers unique insight not seen previously in citation network research.
Limitations.
As our results indicate, the model fits best to communities that are too young to develop a noticeable ageing or attention decay effect. This directly follows from the fact that our model, developed as the minimal analytical extension of the 3DSI model, has no attention decay parameter and thus is not well suited to modelling networks where the attention decay is prominent. To this end, in Appendix A, we propose a variant of our model that includes ageing; while analytically appealing, it unfortunately requires numerical estimation of its parameters. We leave exhaustive examination of its properties as future work. However, the comparison with the decay-equipped version of the model suggests that overcompensation for ageing with increased accidental gain is not present, that is, the estimated values of $1-\rho_{i}$ are not inflated.
Another limitation is that our model inherently relies on the community assignments that promote intra-cluster connection but discourage inter-cluster ones. This is expected in citation networks where communities represent scientific domains, but requires careful consideration in other types of networks. This requirement also disqualifies our model from modelling in-degrees in, e.g., dissociative networks.
Notably, our proposition allows only uniformly distributed citations between communities. While intuitively this assumption may be too strong, as preferential attachment likely happens also between communities, we rely on the accidental terms to capture these relationships in order to reduce the number of parameters of the model. It is possible to derive a similar proposition with a separate preferentiality parameter for each pair of communities, but we have decided otherwise to keep the number of parameters linear with respect to $k$. Relaxing this assumption to allow for preferential attachment between arbitrary pairs of communities would require estimating a $k\times k$ matrix of preferentiality parameters, akin to the SBM model. This would lead to a quadratic explosion in the number of parameters, making the model prone to overfitting and difficult to interpret.
Future work.
We will focus our future works on devising an algorithmic framework for generating benchmark graphs that can be used for community detection, as we believe that our model will be complementary to approaches like SBM or LFR in that it allows for flexible modification of the in-degree sequences inside the communities, allowing for modelling broad families of networks. It would also be beneficial to consider further generalisations of our model in which accidental edges are also distributed non-homogeneously, or in which the parameter values depend on the size of the network as recent research indicates that interdisciplinary citations trends change non-randomly in time Rong et al. (2025).
@@
Appendix A Exponential decay
Let us introduce an alternative formulation of the above model that not only captures the information about the communities, but also assumes exponential attention decay (or ageing) of the nodes. This proposal is directly motivated by the fact that the ageing phenomenon is frequently observed in citation networks and that, as we have seen in the main part of the article, fitting the proposed model to larger communities often suffers from overestimating the degrees of oldest nodes. We assume that both preferential and accidental attractiveness of articles decay exponentially in time and we offer the following explanation: the age is a natural limiting factor for the attractiveness of a research work. Thus, even though the number of citations may be large, its attractiveness will diminish: this is the decay in the preferential term. On the other hand, less popular articles not only gain smaller number of citations preferentially, but due to decreased interest they are less frequent in, for example, search engine results, and are less likely to be cited “out of luck”: this is the decay in the accidental term. For the sake of simplicity and for analytical tractability, we assume the same decay parameter for both types of citations and all communities.
A.1 Preliminary derivation
To formulate the decaying model, we shall first propose an alternative derivation of our non-decaying model based on the following postulates.
  1. 1.
    Preferential growth is linear in expectation, i.e., the expected (multiplicative) growth of a vertex in-degree between times $s$ and $t$ is the ratio of the total numbers citations at times $s$ and $t$. For example, if we assume only preferential growth of the in-degree of vertex $\ell$ between times $s$ and $t$, we have in expectation:
    $d^{(t)}(\ell)=\frac{\Sigma^{(t)}}{\Sigma^{(s)}}d^{(s)}(\ell),$
    where $\Sigma^{(s)}$ and $\Sigma^{(t)}$ are the total citation numbers at times $s$ and $t$.
  2. 2.
    The expected total number of citations gained by community $i$ in local turn $t$ is composed of two elements: community-independent expected accidental gain $\langle a\rangle$, which does not depend on the community (see Eq. (6)), and community-dependent preferential gain $\rho_{i}m_{i}$.
  3. 3.
    Assuming all citations are distributed uniformly between the end of the previous turn and the end of the current turn and that all accidental citations precede all preferential citations (the standard 3DSI assumption), it implies that in turn $t$:
    1. (a)
      first, between local moments t1t-1 and $t-1+\frac{\langle a\rangle}{\langle a\rangle+\rho_{i}m_{i}}$, all $\langle a\rangle$ accidental citations are added,
    2. (b)
      in the remaining time interval from $t-1+\frac{\langle a\rangle}{\langle a\rangle+\rho_{i}m_{i}}$ to $t$, all $\rho_{i}m_{i}$ preferential citations are added.
    Note that this derivation gives a new interpretation to $\nu_{i}=\frac{\rho_{i}m_{i}}{\langle a\rangle+\rho_{i}m_{i}}$. We can treat $1-\nu_{i}$ as the moment in time between the end of turn t1t-1 and the end of turn $t$ when all accidental citations have been added and all preferential citations remain.
  4. 4.
    Following the previous points, we model the growth of the node $\ell$ in community $i$ as a two-step process: additive accidental citation gain and multiplicative preferential citation gain. We thus obtain:
    $d_{i}^{(t)}(\ell)=\left(d_{i}^{(t-1)}(\ell)+\frac{\langle a\rangle}{t-1}\right)\cdot\frac{\Sigma_{i}^{(t)}}{\Sigma_{i}^{(t-1+1-\nu_{i})}}=\left(d_{i}^{(t-1)}(\ell)+\frac{\langle a\rangle}{t-1}\right)\frac{t-1}{t-1-\nu_{i}},$
    which corresponds exactly to the derivation preceding Formula (6).
The above derivation of the standard model allows us to skip calculating preferential gain of individual vertices, as well as exact values of $\Sigma^{(t)}_{i},\Sigma_{i}^{(t-\nu_{i})}$ as long as we can calculate their ratio, which will be useful for composing the recurrence involving decay. More importantly, it allows us to consider the state of the model only at three moments in time: $t-1,t-1+1-\nu_{i}=t-\nu_{i},$ and $t$.
A.2 Decaying model
We assume the same accidental and preferential citation gain as in the non-decaying model, but now the “value” or “mass” of citations taken into account when calculating probabilities is not linearly proportional to the numbers of citations, because we now assume that the citations diminish in the value exponentially over time. We introduce the following notation:
  • $b_{i}^{(t)}(\ell)$ – in-degree of vertex $\ell$ form community $i$ at local time $t$ in the decaying scenario,
  • $q$ – decay parameter such that $0<q\leq 1$; in each turn, existing citations diminish in value times $q$,
  • $Q_{i}^{(t)}$ – total mass of citations in cluster $i$ after the local turn $t$.
For simplicity, we assume that all accidental citations are added exactly at time $t-\nu_{i}$ in turn $t$. Since the initial mass added in each local turn to community $i$ is, without change, $\langle a\rangle+\rho_{i}m_{i}$, and all citations decay with parameter $q$, we have that the total mass satisfies the recurrence:
$Q_{i}^{(t)}=qQ_{i}^{(t-1)}+\langle a\rangle q^{\nu_{i}}+\rho_{i}m_{i}.$
As before, $Q^{(1)}_{i}=0$. The term $q^{\nu_{i}}$ comes from the fact that between adding accidental edges and the end of the turn, $\nu_{i}$ time units elapsed. From this we naturally obtain $Q^{(t)}_{i}=\left(\langle a\rangle q^{\nu_{i}}+\rho_{i}m_{i}\right)[t-1]_{q}$, where we the $q$-number is defined as $n]_{q}=\frac{1-q^{n}}{1-q}$. Please note, however, that since the masses do not grow linearly, the “switch” between accidental and preferential edges does not necessarily happen at $\frac{\langle a\rangle}{\langle a\rangle+\rho_{i}m_{i}}$. We derive this new $\nu_{i}$ from the fact that $Q^{(t-\nu_{i})}_{i}\sim[t-1-\nu_{i}]_{q}$, which leads to:
$\frac{[1-\nu_{i}]_{q}}{[\nu_{i}]_{q}}=\frac{\langle a\rangle}{\rho_{i}m_{i}}\implies\nu_{i}=\log_{q}\left(\frac{\langle a\rangle-\rho_{i}m_{i}+\sqrt{(\langle a\rangle-\rho_{i}m_{i})^{2}+4\langle a\rangle\rho_{i}m_{i}q}}{2\langle a\rangle}\right).$
We do not change the notation for $\nu_{i}$, because we only generalised the previous value. Namely, $\lim_{q\rightarrow 1}\nu_{i}=\frac{\rho_{i}m_{i}}{\langle a\rangle+\rho_{i}m_{i}}$. While deriving the accidental gain in the decaying model of node $\ell$ requires considering global time, we approximate it with its local age. This approach allows us to relate article’s attractiveness only to its age compared to research in the same domain; otherwise, the model would quickly diminish the impact of potentially important papers from slowly growing scientific areas. As such, we assume that the accidental gain of a node $\ell$ is proportional to its weight decaying with age and inversely proportional to the sum of all weights in the same community. We thus write:
$\operatorname*{qAcc}_{i}^{(t)}(\ell)=\frac{\langle a\rangle q^{t-1-l}}{[t-1]_{q}}$
for the accidental decaying gain of node $\ell$ in turn $t$ in community $i$. Gathering all these components together, we formulate the recurrence relation:
$b_{i}^{(t)}(\ell)=\left(b_{i}^{(t-1)}(\ell)q^{1-\nu_{i}}+\operatorname*{qAcc}^{(t)}_{i}(\ell)\right)\frac{Q_{i}^{(t)}}{Q_{i}^{(t-\nu_{i})}}=\left(b_{i}^{(t-1)}(\ell)q^{1-\nu_{i}}+\frac{\langle a\rangle q^{t-1-l}}{[t-1]_{q}}\right)\frac{[t-1]_{q}}{[t-1-\nu_{i}]_{q}}.$
The factor $q^{1-\nu_{i}}$ comes from the fact that if we assume all accidental citations are added at moment $t-\nu_{i}$, the previous citations have decayed over the interval $1-\nu_{i}$. Following a similar telescopic process for unfolding the recurrence, we obtain the closed-form formula:
$b_{i}^{(t)}(\ell)=\frac{\langle a\rangle q^{t+\nu_{i}-\ell-1}}{[\nu_{i}]_{q}}\left(q^{\nu_{i}(\ell-t)}\frac{\Gamma_{q}(t)\Gamma_{q}(\ell-\nu_{i})}{\Gamma_{q}(t-\nu_{i})\Gamma_{q}(\ell)}-1\right),$
Figure 5: Decaying model-predicted in-degree sequences for the same graph as presented in Figure 1, this time with varying $q$.
where $\Gamma_{q}$ is the $q$-analog of the Gamma function. Naturally we have $\lim_{q\rightarrow 1}b_{i}^{(t)}(\ell)=d_{i}^{(t)}(\ell)$. Figure 5 presents how varying $q$ impacts the resulting in-degree sequence in different scenarios. Even though the analytical form of the decaying model is visually similar to the non-decaying model, gamma q-analogs are much more difficult to handle both analytically and numerically. The decaying model currently suffers from the following limitations.
  • Most parameters need to be estimated numerically; one disadvantage of exponential decay is that the decay parameter $q$ will have values near $1$ for all reasonably-sized networks, which makes the numerical estimation even harder.
  • The analytical form contains a number of values that either grow or diminish exponentially. This causes the decaying model to be prone to numerical instabilities both during estimation and plotting.
  • While $Q^{(t)}_{i}$ correctly models the effective attractiveness of publications in community $i$, it loses its interpretability as the number of citations gained. Specifically, unless $q$ is carefully estimated, the model could produce in-degree sequences that are non-monotonic with respect to age, since this model allows – and even requires – “losing citations”.
For completeness, we show how this decaying model with numerically estimated parameters fits to the two previously considered real networks with exogenous labels in Figures 6 and 7.
Figure 6: Complementary cumulative distribution functions of the observed and modelled in-degree sequences in the seven clusters of the DBLP authors dataset. The numerical parameter estimation was done using L-BFGS-B with mean squared logarithmic error loss. The last plot shows the full in-degree sequences.
Figure 7: Complementary cumulative distribution functions of the observed and modelled in-degree sequences in the seven clusters of the Cora dataset. The numerical parameter estimation was done using L-BFGS-B with mean squared logarithmic error loss. The last plot shows the full in-degree sequences. The numerical non-decaying distribution overlaps nearly everywhere with the decaying distribution.
For numerical parameter estimation, we minimised the mean squared logarithmic error using the L-BFGS-B algorithm Zhu et al. (1997) with 100 restarts. To level the field, we fitted the standard version of the model (ND=non-decaying) both numerically as well as using the previously derived estimators. To resolve ambiguous optimality of $\rho_{i}m_{i}$, we required $m_{i}\sim\Sigma^{(t)}_{i}$ in the non-decaying model and $m_{i}\sim\frac{Q^{(t)}_{i}}{[t-1]_{q}}$ in the decaying model. We observe that while the fit to the tails has mostly improved over the analytical non-decaying model, it has often happened at the cost of worsening the fit to a single community – Mathematical Optimisation in DBLP, and Theory in Cora network, respectively. These artefacts may also stem from the difficulty posed by numerically fitting the decaying model due to high sensitivity of the $q$ parameter. At this point, we observe limited advantage of introducing the exponential decay into the model, albeit with proper numerical estimation it may produce better fits to the data over the non-decaying model. We leave these potential improvements, as well as considering other decay modelling approaches, as future work.
Table 3: Estimated values of the model parameters for all communities in both networks derived from the numerically fitted decaying model.
Cora (q=0.999999898458) DBLP authors (q=0.999999928008)
Cluster $\hat{m_{i}}$ $\hat{\rho_{i}}$ Cluster $\hat{m_{i}}$ $\hat{\rho_{i}}$
Case Based $1.919$ $0.308$ Computer network $45.206$ $0.822$
Genetic Algorithms $2.179$ $0.749$ Computer science $95.130$ $0.546$
Neural Networks $1.749$ $0.472$ Computer vision $67.503$ $0.896$
Probabilistic Methods $1.893$ $0.409$ Control theory $30.129$ $0.903$
Reinforcement Learning $2.274$ $0.583$ Electronic engineering $23.059$ $0.766$
Rule Learning $1.902$ $0.395$ Math. optimisation $44.415$ $0.906$
Theory $2.107$ $0.405$ Mathematics $29.555$ $0.999$
Pattern recognition $77.382$ $0.656$
Table 3 presents the numerically estimated parameters on both datasets in the decayed model. We observe that to alleviate the decay, most estimates of $\rho_{i}$ in the Cora network are lower than in Table 2.